OpenStudy (beth12345):

What would be the gravitational field strength 2000km above the Earth? 50km above the Earth?

OpenStudy (anonymous):

Gravity decreases with altitude, since greater altitude means greater distance from the Earth's center. so the gravity at 2000 km above the earth surface will be very less as compare to 50 km above the earth surface.

OpenStudy (anonymous):

\[F=GmM/(R _{e}+r)^2\] where M is mass of the Earth & Re is radius of the Earth

OpenStudy (beth12345):

what is Gm and r?

OpenStudy (anonymous):

r is distance of mass to Earth and G is gravitation constant that equal to 6.63*10^34(i'm not sure about G's value but i think that.many of physics book has this eg holliday)

OpenStudy (anonymous):

G is the gravitational constant and m is the mass of any object on the earth surface it is not r it is h which means

OpenStudy (anonymous):

i think hosein it is h because the question stats that is about the altitude

OpenStudy (beth12345):

uhm, I don't really know the numbers to plug into the equation

OpenStudy (anonymous):

yea here r is h

OpenStudy (anonymous):

you can't solve this,bath?

OpenStudy (anonymous):

so u can put 2000km as h or 5o as h but u cannot solve it

OpenStudy (beth12345):

ya

OpenStudy (beth12345):

i don't know what m is

OpenStudy (anonymous):

m is mass of 2nd body

OpenStudy (anonymous):

m is the mass of a any object on the earth surface or the object above the earth surface e.g a satellite

OpenStudy (beth12345):

do you know what value I would plug into it?

OpenStudy (anonymous):

mass of an object e.g 40 kg but is negligible as compared to earths mass

OpenStudy (anonymous):

G=6.67*10(-11) N(m/kg)^2 M = 5.9722 × 1024 kg

OpenStudy (anonymous):

subtitute the values given if u have any word problem

OpenStudy (anonymous):

you ought be consider that in formula

OpenStudy (beth12345):

ok so would the equation then become \[G= 6.67\times10^{-11}\] \[M= 5.98 \times10^{24}\] \[R e = 6.38 \times10^{6 }\] \[h= (2000)or (50)\]

OpenStudy (anonymous):

yea put it in formula

OpenStudy (anonymous):

you should these standard values in the formula

OpenStudy (beth12345):

and then \[((6.67\times10^{-11}) \times 40) \div ((6.38 \times 10^{6}) + 2000)\]

OpenStudy (anonymous):

where is Earth mass? and denominator has power 2

OpenStudy (anonymous):

but accelaration due to gravity too comes in the formula

OpenStudy (anonymous):

we forgot the main thing

OpenStudy (anonymous):

what thing?sgholap100

OpenStudy (anonymous):

accelaration due to gravity

OpenStudy (anonymous):

ican't understand your goal?can you explane

OpenStudy (anonymous):

you can use gRe^2 instead of GMe

OpenStudy (anonymous):

as the question say that the object is above earth surface accelaration due to gravity(g=9.8m/s^2) too comes then the formula becomes

OpenStudy (anonymous):

your goal is?\[mg=GMm/R _{e}^2\] am i right?

OpenStudy (beth12345):

\[((6.67 \times 10 ^{-11}) \times 40) \times 5.98 \times 10 ^{24}) \div(6.38 \times 10 ^{6} +2000)^{2} \]

OpenStudy (anonymous):

oh yeah i am sorry its was my mistake i was considering Effect of altitude on gravity @hosein u r right

OpenStudy (anonymous):

ithink it's right.

OpenStudy (anonymous):

yes you r right

OpenStudy (beth12345):

so would the answer be 391.7 ?

OpenStudy (beth12345):

ok ya I think I got it, thanks guys :)

OpenStudy (anonymous):

welcome but do one thing neglect the m and h in the formula as they are very small as compared to M and R OK ^_^